OpenStudy (alex_mattucci):

Please Help Me!!! Need help with some polynomial problems. Marvel Man movie is being released to a limited number of movie theaters. The number of showtimes for the movie is twice the number of movie theaters. The number of seats available for each showtime is three times the number of showtimes. Let x represent the number of theaters showing the movie. A. Determine the three factors that are to be multiplied to determine the total number of seats available for the movie. B. When simplified what is the degree and classification of this expression? C. Interpret each pairing of factors as a single unit in the context.

OpenStudy (anonymous):

hmmm..

OpenStudy (anonymous):

im leaning agansed a, or b.

OpenStudy (alex_mattucci):

@SolomonZelman and @saifoo.khan Please Help Me!!!

OpenStudy (candyme):

M=(x)(s)(c) Movie Seats = (#of theaters)(#of shows)(# number of seats) s=2x c=3s so you want it all in terms of x start replacing each thing replace s with 2x M=(x)(2x)(3s) Now figure out what c is if s=2x, then c=3(2x)=6x Your final answer is M=(x)(2x)(6x) This keeps everything in terms of one variable

OpenStudy (candyme):

Now simplify

OpenStudy (candyme):

\[M=(12x^2)(x)\]

OpenStudy (candyme):

\[M=12x^3\]

OpenStudy (candyme):

Since there is only one little piece in the equation now, It is a monomial. The degree is the highest exponent, which is 3.

OpenStudy (candyme):

Does this help?

OpenStudy (alex_mattucci):

Yes thank you!